OpenStudy (anonymous):

What's the ordered pair for y=6x+5 y=-7x+5

OpenStudy (radar):

y=y 6x+5=-7x+5 13x=0 x=0/13=0 y=5 (0,5)

OpenStudy (anonymous):

Thanks=)